# Reconstructing People, Places, and Cameras

### v1

@@ Abstract
We present “Humans and Structure from Motion” (HSfM), a method for jointly reconstructing multiple human meshes, scene point clouds, and camera parameters in a metric world coordinate system from a sparse set of uncalibrated multi-view images featuring people. Our approach combines data-driven scene reconstruction with the traditional Structure-from-Motion (SfM) framework to achieve more accurate scene reconstruction and camera estimation while simultaneously recovering human meshes. In contrast to existing scene reconstruction and SfM methods that lack metric scale information, our method estimates approximate metric scale by leveraging the human statistical model. Furthermore, our method reconstructs multiple human meshes within the same world coordinate system with the scene point cloud, effectively capturing spatial relationships among individuals and their positions in the environment. We initialize the reconstruction of humans, scenes, and cameras using robust foundational models and jointly optimize these elements. This joint optimization synergistically improves the accuracy of each component. We compare our method with existing methods on two challenging benchmarks, EgoHumans and EgoExo4D, demonstrating significant improvements in human localization accuracy within the world coordinate frame (reducing error from 3.51m to 1.04m in EgoHumans and from 2.9m to 0.56m in EgoExo4D). Notably, our results show that incorporating human data into the SfM pipeline improves camera pose estimation (e.g., increasing RRA@15 by 20.3% on EgoHumans). Additionally, qualitative results show that our approach improves scene reconstruction quality. Our code is available at muelea.github.io/hsfm.
@@ 1 Introduction
In summary, we present Humans and Structure from Motion (HSfM), which provides a comprehensive representation of the world—encompassing people, places, and cameras—marking a step forward in understanding complex real world environments. Please find more results at the project website: muelea.github.io/hsfm.
@@ 2 Related Work
Table 1: Comparison of methods across different features. Previous works in human pose estimation like HMR2, SLAHMR, Multi-HMR, and UnCaliPose has made great progress in reconstructing body poses in single- and multi-person setups from images. Recent methods like DUSt3R and MASt3R are able to recover accurate camera poses and scene points clouds (places). This includes reconstructions with statistically correct scale (Stat. Scale) which can be obtained from human body models, e.g. in SLAHMR, or, as in MASt3R, from world knowledge. Our approach, HSfM, is the first method that can jointly reconstruct multiple people, scene, and cameras from sparse uncalibrated multi-view images.
@@
Structure from Motion. Structure from Motion (SfM) [5, 15, 19] aims to reconstruct camera poses and 3D scene geometry from a set of images by establishing pixel correspondences across views. Traditional SfM Methods [1, 9, 10], such as COLMAP [41, 42] employ keypoint detection, matching based on locally invariant descriptors [30, 4, 40], and incremental bundle adjustment to estimate camera poses and sparse 3D points. However, these traditional approaches are highly sensitive to noise at each stage of their sequential pipeline and require specific conditions for input, such as a large number of camera views with substantial overlapping image areas.
@@ 4 Humans and Structure from Motion
Our method takes as input an uncalibrated, sparse set of images capturing people in a scene at a single moment in time. Given this input, our goal is to jointly estimate each person’s human parameters, the scene, and the camera parameters. Our key insight is that jointly reasoning about people, scene structure, and cameras improves all three aspects of reconstruction. To achieve this, we integrate the global scene optimization approach from recent scene reconstruction methods with the traditional Structure-from-Motion (SfM) formulation. This integration leverages 2D human keypoints as reliable correspondences and 3D human meshes as robust 3D structures for bundle adjustment. Please also refer to Figure 2.
Our joint optimization approach offers several key advantages. By incorporating human mesh predictions, the method introduces metric scale into the reconstruction process, leveraging statistical information about human body dimensions. The cameras and 2D human keypoints enable precise positioning of individuals within the world coordinate system, allowing for the recovery of their heights and relative distances. Additionally, correspondences of 2D human keypoints enhance camera calibration, which in turn improves scene reconstruction. The scene structure further stabilizes the camera pose registration, creating a feedback loop that refines the overall system. The result is a globally consistent reconstruction of humans, scenes, and cameras, providing a comprehensive understanding of the environment.
Since this is an under-determined problem, we take advantage of data-driven 3D human and scene reconstruction methods [16, 53] to provide initializations. In the following sections, we describe how we align these independent entities into a common world coordinate frame, followed by the joint optimization of the people, scene, and camera.
@@ 4.1 Initialization of World
One may simply start the optimization by setting the scale $\alpha=1$. However, since SfM reconstructions are up-to-scale, the magnitude of $t$ may vary significantly in every reconstruction. If the SfM scene is too small relative to people, cameras may be in front of the humans, e.g. when the scene is placed inside the human mesh, leading to degenerate solutions. Setting alpha to an arbitrarily big value such as $\alpha=100$ can prevent these issues, but makes the problem prone to local minima.
We propose an analytical method to approximate a consistent reconstruction using data-driven outputs of 2D/3D human keypoints and camera parameters. Initially, we roughly position individuals within the scene by estimating $\gamma$, based on the predicted 2D/3D human keypoints and the focal length from $\tilde{K}$, following a similar approach to Ye et al. [61]. Next, we calculate the initial scale $\alpha$ of the data-driven SfM-predicted world by aligning the data-driven camera positions with human-centric camera positions.
@@
Finally, given the camera positions $\hat{T}$ derived from the humans and the data-driven camera position predictions $\tilde{T}$, we solve a least-squares problem to compute $\hat{\alpha}$, which aligns the scene pointmap $\mathcal{S}$ and the camera translation $t$ with the metric-scale world defined by the humans $\tilde{H}$. The scaling factor $\hat{\alpha}$ provides a reasonable approximation of the metric scale. For the initial estimates of $\tilde{H}$, we rely on an estimate from a reference camera in practice. Please refer to the supplementary video on our project page for an animated illustration of this approach.
@@ 4.2 Reconstructing People, Places, and Cameras
Global scene optimization: We adapt the global alignment loss from DUSt3R [53], which originally optimizes camera and world pointmaps.
Intuitively, this loss takes pairs of cameras, say Alice and Bob ($c_{i}$ and $c_{j}$), with the predicted content of Alice in Bob’s view. The alignment loss transforms this “Alice’-content-in-Bob’s-view” to the world coordinate frame where it’s compared against the optimized global scene pointmap, which is Alice’s content in the world.
@@
Figure 3: Qualitative results from HSfM. We show our optimized result on sequences from EgoHumans (top) and EgoExo4D (bottom). Note how in the Initial state (left) people are floating in the air (a), how the scene and human scale is not aligned (e), and how noisy the scene appears (c). Our method resolves these problems by grounding people in the scene (b), recovering plausible metric scale (f), and better camera estimates (d). We achieve this with minimal assumptions and without scene contact constraints.
Figure 4: Qualitative results in the wild. We show reconstructions on in-the-wild images taken with two smartphones (a), demonstrating the reconstruction of humans and scenes. Unlike previous works [70, 61], which adopt human-scene contact priors that hinder generalization to scenarios without ground foot contact, HSfM recovers accurate world locations of the human meshes that are coherent with the static scene structure. The use of humans in our framework (c) not only serves as a reliable initialization for 3D structure in the SfM formulation but also provides more faithful and complete information about people in the world, which a noisy human point cloud (b) cannot offer. For visualization purposes, the human point cloud is removed using SAM2 [38].
@@ 5 Experiments
EgoExo4D UnCaliPose* [55] 2.90 - 0.14 2.43 1.16 65.61 0.19 0.29 - - 0.24 0.34
DUSt3R [53] - - - - 0.33 9.92 0.81 0.88 - - 0.64 0.85
MASt3R [26] - - - 0.96 0.35 11.70 0.79 0.87 0.06 0.14 0.68 0.81
HSfM (init.) 5.29 - 0.07 1.27 0.33 9.92 0.81 0.88 0.05 0.09 0.64 0.85
HSfM (Ours) 0.56 - 0.06 0.95 0.36 11.57 0.78 0.87 0.07 0.16 0.67 0.82
Table 2: Evaluation on EgoHumans and EgoExo4D. HSfM outperforms existing human and scene reconstruction methods, delivering metric-scale reconstructions for humans, the scene, and cameras within the same world coordinate system. Our approach shows major improvement over the initial estimates HSfM (init.), obtained from DUSt3R [53] and HMR2 [16] (Section 4.1), particularly when multiple humans are present in the scene, as seen in EgoHumans, compared to a single individual in EgoExo4D. Additionally, HSfM surpasses MASt3R in metric-scale camera metrics, demonstrating the benefit of incorporating humans into the Structure-from-Motion (SfM) pipeline for scale recovery. All baselines and our method use four cameras as input on EgoHumans and up to six cameras on EgoExo4D. Human metrics and camera translation metrics are reported in meters.
@@
We evaluate the effectiveness of our approach in terms of human pose estimation within the world coordinate system and camera accuracy. Additionally, we present qualitative results showcasing the effects of our joint optimization on humans, scene pointmaps, and cameras.
5.1 Evaluation
Evaluation Datasets: We evaluate on the EgoHumans [24] and EgoExo4D [17] datasets. EgoHumans is a new multi-view, multi-human video benchmark designed for egocentric 3D human pose estimation and tracking. It captures over 125,000 images of two to four people interacting in dynamic, real-world activities. EgoExo4D is a large-scale, multi-view video dataset comprising simultaneously captured egocentric and exocentric videos of skilled human activities such as sports, music, dance, and bike repair for a single person. Please refer to the supplementary material for more details.
@@ 5.2 Results
We report the results in Tab. 2, comparing various human and camera estimation baselines on EgoHumans and EgoExo4D. We evaluate human and camera estimates against UnCaliPose [55], and camera estimates against DUSt3R [53] and MASt3R [26]. UnCaliPose is the closest baseline, as it uses humans for Structure-from-Motion (SfM) and jointly reconstructs both humans and cameras. However, it does not reconstruct the scene, and the scale is recovered using ground truth bone lengths during testing. For a fair comparison, we assume known re-identification across views and use ViTPose [59] for both UnCaliPose and HSfM. Additionally, we use the predicted focal length from DUSt3R [53] for UnCaliPose. Since no method exists that jointly estimates people, places, and cameras at metric scale from sparse multi-view images, we further report results on HSfM (init.), i.e. the world state after the initialization procedure described in Sec. 4.1.
@@
As shown in the EgoHumans table, HSfM (init.) outperforms UnCaliPose in scale-normalized human metrics, reducing GA-MPJPE by approximately 24% and PA-MPJPE by over 50%. Additionally, HSfM (init.) achieves a relative rotation accuracy that is nearly double, with RRA@10 improving by 86% and RRA@15 by over 100% compared to UnCaliPose. This demonstrates the effectiveness of our strong initialization, which leverages foundational models—HMR2 for humans and DUSt3R for cameras.
@@
Our optimization yields substantial improvements. For instance, W-MPJPE is reduced from 4.28m in HSfM (init.) to 1.04m in HSfM, demonstrating the effectiveness of our approach in resolving scale ambiguity and accurately positioning humans within the world. Additionally, the camera metrics show significant enhancements. Our method surpasses the initial outputs of DUSt3R and outperforms MASt3R in metric-scale camera metrics with large margins, reducing TE from 4.97m to 2.09m and achieving approximately seven times better CCA@15. These results highlight the advantages of incorporating humans into the reconstruction process to achieve a consistent metric-scale world, consistent with the findings of Zhao et al. [67].
Similar trends are observed in the EgoExo4D results, further validating the effectiveness of our method in reconstructing the world. HSfM achieves substantially better human metrics compared to UnCaliPose, reducing W-MPJPE from 2.90m to 0.56m and PA-MPJPE from 0.13m to 0.06m. For metric-scale camera metrics, HSfM outperforms both MASt3R and HSfM (init.), improving CCA@15 by 14% and 78%, respectively.
However, the improvements, especially in scale-invariant camera metrics (e.g., RRA@10/15) are less compared to those observed in EgoHumans. This reduced performance is likely due to two factors: significant occlusions in indoor images, which affect 2D human keypoint predictions, and the fact that only a single person is used for optimization in EgoExo4D. In contrast, EgoHumans utilizes multiple individuals, providing a greater number of 2D human keypoint correspondences, which strengthens the optimization process. Our ablation study confirms this trend.
@@
Ablations: We first validate the importance of our joint optimization of humans, scenes, and cameras in Table 3. The first variant, M1, detaches the gradients from the human loss to the scene and camera parameters while still optimizing all parameters. Essentially, the cameras and scene do not adjust to minimize the human losses. This leads to W-MPJPE improving only marginally from the initialization (4.28m to 3.94m) and GA-MPJPE slightly increasing to 0.57m from 0.51m. Since human losses do not influence the optimization of the camera and scene, the camera metrics also stagnate, with RRA@15 showing no improvement, and CCA@15 increasing only marginally by 5.3%. The second variant, M2, optimizes cameras and humans solely based on the human loss, excluding the scene loss. Interestingly, this significantly improves W-MPJPE from 4.28m to 1.29m and GA-MPJPE from 0.57m to 0.24m, indicating accurate recovery of human world locations and relative distances. However, the camera metrics degrade considerably: CCA@15 decreases by 36.8%, and RRA@15 declines by 7.6%. This highlights that without incorporating scene losses, the structure fails to anchor the cameras, leading to overfitting to human keypoints. This behavior is similar to the limitations observed in UnCaliPose, which relies solely on human keypoints for Structure-from-Motion (SfM). In contrast, our full method (M3: HSfM (Ours)) achieves the best performance across all metrics, reducing W-MPJPE to 1.04m, GA-MPJPE to 0.21m, and improving camera metrics, with RRA@15 increasing by 22% and CCA@15 by 21%. These results underscore the critical importance of jointly optimizing humans, scenes, and cameras to achieve a coherent and accurate metric-scale reconstruction.
@@
The results highlight that increasing the number of humans introduces more correspondences for Structure-from-Motion (SfM), thereby strengthening the bundle adjustment process. This validates our strategy of integrating global scene optimization techniques from recent scene reconstruction methods with the traditional SfM formulation. Our approach effectively leverages 2D human keypoints as reliable correspondences and 3D human meshes as robust structures to enhance bundle adjustment and produce coherent, metric-scale reconstructions. Please refer to the supplementary material for the ablation study on camera/scene scale initialization and the impact of the number of cameras.
Qualitative Results: See Figure 1 and Figure 3 for qualitative results on EgoHumans and EgoExo4D, and Figure 4 for reconstructions on images taken in the wild.
Figure 3 shows the intermediate steps of the optimization, where in the beginning, people are floating around in mid air. After the optimization, their feet are consistent with the environment, without any explicit contact constraints. We see that the structure also improves — the pointmaps are more coherent around the ground.
Lastly, in Figure 4, we present visualizations of HSfM’s outputs on in-the-wild images. The data was captured using a minimal setup consisting of two cell phones, two tripods, and the Riverside app [^1] https://riverside.fm/ for straightforward time synchronization. Despite this simple setup, our multi-view optimization algorithm successfully handles challenging scenes, such as individuals jumping, without relying on any heuristic contact priors and small data-driven motion priors that previous works [61, 44, 68] use. Please refer to the supplementary material for more qualitative results.

### v2

@@ Abstract
We present “Humans and Structure from Motion” (HSfM), a method for jointly reconstructing multiple human meshes, scene point clouds, and camera parameters in a metric world coordinate system from a sparse set of uncalibrated multi-view images featuring people. Our approach combines data-driven scene reconstruction with the traditional Structure-from-Motion (SfM) framework to achieve more accurate scene reconstruction and camera estimation while simultaneously recovering human meshes. In contrast to existing scene reconstruction and SfM methods that lack metric scale information, our method estimates approximate metric scale by leveraging the human statistical model. Furthermore, our method reconstructs multiple human meshes within the same world coordinate system with the scene point cloud, effectively capturing spatial relationships among individuals and their positions in the environment. We initialize the reconstruction of humans, scenes, and cameras using robust foundational models and jointly optimize these elements. This joint optimization synergistically improves the accuracy of each component. We compare our method with existing methods on two challenging benchmarks, EgoHumans and EgoExo4D, demonstrating significant improvements in human localization accuracy within the world coordinate frame (reducing error from 3.59m to 1.04m in EgoHumans and from 3.01m to 0.50m in EgoExo4D). Notably, our results show that incorporating human data into the SfM pipeline improves camera pose estimation (e.g., increasing RRA@15 by 20.3% on EgoHumans). Additionally, qualitative results show that our approach improves scene reconstruction quality. Our code is available at muelea.github.io/hsfm.
@@ 1 Introduction
In summary, we present Humans and Structure from Motion (HSfM), which provides a comprehensive representation of the world—encompassing people, places, and cameras—marking a step forward in understanding complex real world environments.
@@ 2 Related Work
Table 1: Comparison of methods across different features. Previous works in human pose estimation like HMR2, SLAHMR, Multi-HMR, and UnCaliPose has made great progress in reconstructing body poses in single- and multi-person setups from images. Recent methods like DUSt3R and MASt3R are able to recover accurate camera poses and scene points clouds (places). This includes reconstructions with statistically correct scale (Stat. Scale) which can be obtained from human body models, e.g. in SLAHMR, or, as in MASt3R, from world knowledge. Our approach, HSfM, is the first to jointly reconstruct multiple people, scene, and cameras from sparse uncalibrated multi-view images.
@@
Structure from Motion. Structure from Motion (SfM) [5, 15, 19] aims to reconstruct camera poses and 3D scene geometry from a set of images by establishing pixel correspondences across views. Traditional SfM Methods [1, 9, 10], such as COLMAP [41, 42] employs keypoint detection, matching based on locally invariant descriptors [30, 4, 40], and incremental bundle adjustment to estimate camera poses and sparse 3D points. However, these traditional approaches are highly sensitive to noise at each stage of their sequential pipeline and require specific conditions for input, such as a large number of camera views with substantial overlapping image areas.
@@ 4 Humans and Structure from Motion
Our method takes as input an uncalibrated, sparse set of images capturing people in a scene at a single moment in time. Given this input, our goal is to jointly estimate each person’s human parameters, the scene, and the camera parameters. Our key insight is that jointly reasoning about people, scene structure, and cameras improves all three aspects of reconstruction. To achieve this, we integrate global scene optimization from recent scene reconstruction methods with the traditional Structure-from-Motion (SfM) formulation. This integration leverages 2D human keypoints as reliable correspondences and 3D human meshes as robust 3D structures for bundle adjustment. Please also refer to Figure 2.
Our joint optimization approach has several advantages. By incorporating human mesh predictions, the method introduces metric scale into the reconstruction process, leveraging statistical information about human body dimensions. The cameras and 2D human keypoints enable precise positioning of individuals within the world coordinate system, allowing for the recovery of their heights and relative distances. Additionally, correspondences of 2D human keypoints enhance camera calibration, which in turn improves scene reconstruction. The scene structure further stabilizes the camera pose registration, creating a feedback loop that refines the overall system. The result is a globally consistent reconstruction of humans, scenes, and cameras, providing a comprehensive understanding of the environment. Since this is an under-determined problem, we take advantage of data-driven 3D human [16] and scene [53] reconstruction methods to provide initializations. Note that our method can easily integrate other mesh regressors tailored towards estimating non-standard body size such as BEV [46].
@@ 4.1 Initialization of World
One may simply start the optimization by setting the scale $\alpha=1$. However, since SfM reconstructions are up-to-scale, the magnitude of $t$ may vary significantly in every reconstruction. If the SfM scene is too small relative to people, cameras may be in-front of the humans, e.g. when the scene is placed inside the human mesh, leading to degenerate solutions. Setting alpha to arbitrary big value can prevent this, but makes the problem prone to local minima.
We propose an analytical method to approximate a consistent reconstruction using data-driven outputs of 2D/3D human keypoints and camera parameters. Initially, we roughly position individuals within the scene by estimating $\gamma$, based on the predicted 2D/3D human keypoints and the focal length from $\tilde{K}$, following a similar approach to Ye et al. [61]. Next, we calculate the initial scale $\alpha$ of the data-driven SfM-predicted world by aligning the data-driven with human-centric camera positions.
@@
Finally, given the camera positions $\hat{T}$ derived from the humans and the data-driven camera position predictions $\tilde{T}$, we solve a least-squares problem to compute $\hat{\alpha}$, which aligns the scene pointmap $\mathcal{S}$ and the camera translation $t$ with the metric-scale world defined by the humans $\tilde{H}$. The scaling factor $\hat{\alpha}$ provides a reasonable approximation of the metric scale. For the initial estimates of $\tilde{H}$, we rely on an estimate from a reference camera in practice. Please see the project page supplementary video for an animated illustration.
@@ 4.2 Reconstructing People, Places, and Cameras
Global scene optimization: We adapt the global alignment loss from DUSt3R [53], which originally optimizes camera and world pointmaps. Intuitively, this loss takes pairs of cameras, say A and B ($c_{i}$ and $c_{j}$), with the predicted content of $c_{i}$ in B’s view. The alignment loss transforms this “A’s-content-in-B’s-view” to the world coordinate frame where it’s compared against the optimized global scene pointmap, which is A’s content in the world.
@@
Figure 3: Qualitative results from HSfM. We show our optimized result on sequences from EgoHumans (top) and EgoExo4D (bottom). Note how in the Initial state (left) people are floating in the air (a), how the scene and human scale is not aligned (e), and how noisy the scene appears (c). Our method resolves these problems by grounding people in the scene (b), recovering plausible metric scale (f), and better camera estimates (d). We achieve this without scene contact constraints, which often require assumptions about the environment—such as flat terrain—or about motion, such as the assumption that humans are always in contact with the ground (i.e., no jumping). For more qualitative results, including a demo on images taken in the wild with a minimal capturing setup, please see our supplementary material.
@@ 5 Experiments
EgoExo4D UnCaliPose* [55] 3.59 - 1.19 2.21 0.98 63.98 0.20 0.31 - - 0.26 0.37
DUSt3R [53] - - - - 0.34 10.06 0.81 0.88 - - 0.64 0.84
MASt3R [26] - - - 1.03 0.36 9.11 0.81 0.90 0.09 0.17 0.70 0.81
HSfM (init.) 5.80 - 0.08 1.27 0.34 10.06 0.81 0.88 0.05 0.10 0.64 0.84
HSfM (Ours) 0.50 - 0.07 1.01 0.34 10.39 0.80 0.89 0.05 0.14 0.70 0.84
Table 2: Evaluation on EgoHumans and EgoExo4D. HSfM outperforms existing human and scene reconstruction methods, delivering metric-scale reconstructions for humans, the scene, and cameras within the same world coordinate system. Our approach shows major improvement over the initial estimates HSfM (init.), obtained from DUSt3R [53] and HMR2 [16] (Section 4.1), particularly when multiple humans are present in the scene, as seen in EgoHumans, compared to a single individual in EgoExo4D. Additionally, HSfM surpasses MASt3R in metric-scale camera metrics, demonstrating the benefit of human size for recovering scene scale. All baselines and our method use four cameras on EgoHumans and up to six cameras on EgoExo4D. Human and camera translation metrics are reported in meters.
@@
We evaluate HSfM’s effectiveness in terms of human pose estimation within the world coordinate system and camera accuracy, and show qualitative results of our joint optimization on humans, scene pointmaps, and cameras.
Evaluation Datasets: We evaluate on EgoHumans [24] and EgoExo4D [17]. EgoHumans is a multi-view, multi-human benchmark for human pose estimation, featuring videos of 2-4 people in real-world activities. EgoExo4D is a large-scale dataset of people performing tasks like dancing, playing music, or bike repair; see Sup. Mat. for details.
@@ 5.1 Results
We compare human and camera estimation baselines on EgoHumans and EgoExo4D (Tab. 2), including UnCaliPose [55], DUSt3R [53], and MASt3R [26]. UnCaliPose jointly reconstructs humans and cameras using SfM but does not reconstruct the scene and relies on ground truth bone lengths during testing. For a fair comparison, we assume known re-identification across views, use ViTPose for UnCaliPose and HSfM, and apply DUSt3R’s focal length to UnCaliPose. Since no existing approach jointly estimates people, places, and cameras at metric scale from sparse multi-view images, we also report HSfM (init.), the state after our initialization in Sec. 4.1.
@@
On EgoHumans, HSfM (init.) outperforms UnCaliPose in scale-normalized human metrics, reducing GA-MPJPE by approximately 24% and PA-MPJPE by over 50%. It nearly doubles relative rotation accuracy compared to UnCaliPose with RRA@10 and RRA@15 improvements of 86% and 100%, respectively. These gains show the strength of leveraging HMR2 and DUSt3R for initialization.
@@
Our optimization further improves results: W-MPJPE drops from 4.28m in HSfM (init.) to 1.04m in HSfM, demonstrating the effectiveness of our approach in resolving scale ambiguity and accurately positioning humans within the world. The camera metrics also improve substantially, surpassing DUSt3R’s initial outputs and outperforming MASt3R, reducing TE from 4.97m to 2.09m and achieving about seven times better CCA@15. These results highlight the advantages of incorporating humans into the reconstruction process to achieve a consistent metric-scale world, consistent with the findings of Zhao et al. [68].
Similar trends observed in EgoExo4D further validate the effectiveness of our method. HSfM achieves substantially better human metrics compared to UnCaliPose, reducing W-MPJPE from 3.59m to 0.50m and PA-MPJPE from 0.13m to 0.07m. For metric-scale camera metrics, HSfM outperforms HSfM (init.), improving CCA@15 by 33%. The improvements in scale-invariant camera metrics are smaller than on EgoHumans, likely due to heavy indoor occlusions affecting 2D keypoint predictions and the use of only one person in EgoExo4D. Similarly, MASt3R estimates slightly more accurate camera centers (CCA@10/15) on EgoExo4D while HSfM is on-par and slightly better on the scale-invariant version (s-CCA@10/15). This is likely due to a single person being less effective for estimating scene scale; an effect we ablate in Tab. 3 on EgoHumans where we also observe a single person to be less efficient for estimating scene scale compared to more people. In contrast, EgoHumans benefits from multiple individuals, providing more 2D keypoint correspondences and strengthening optimization. Our ablation study confirms this trend.
@@
Ablations: We validate the importance of jointly optimizing humans, scenes, and cameras in Table 3. The first variant, M1, detaches the gradients from the human loss to the scene and camera parameters while still optimizing all parameters. Essentially, the cameras and scene do not adjust to minimize the human losses. This leads to minor W-MPJPE improvement (4.28m to 3.94m) and slightly higher GA-MPJPE (0.51m to 0.57m). Camera metrics nearly stagnate (RRA@15, CCA@15) since human losses do not influence the optimization of camera and scene. The second variant, M2, optimizes cameras and humans solely based on the human loss, excluding the scene loss. Interestingly, this significantly improves W-MPJPE (4.28m to 1.29m) and GA-MPJPE (0.57m to 0.24m), indicating accurate recovery of human world locations and relative distances. However, the camera metrics degrade considerably: CCA@15 by 36.8%, and RRA@15 by 7.6%. Without scene losses, the structure fails to anchor the cameras and overfits to human keypoints. This behavior is similar to the limitations observed in UnCaliPose, which relies solely on human keypoints for SfM. In contrast, our full method (M3) achieves the best metrics, reducing W-MPJPE and GA-MPJPE to 1.04m and 0.21m, respectively, and increasing RRA@15 by 22% and CCA@15 by 21%. This highlights the importance of jointly optimizing humans, scenes, and cameras to achieve a coherent and accurate metric-scale reconstruction.
@@
The results highlight that increasing the number of humans, i.e. introducing more correspondences for Structure-from-Motion (SfM), strengthens the bundle adjustment process. This validates our strategy of integrating global scene optimization techniques from recent scene reconstruction methods with the traditional SfM formulation. Our approach effectively leverages 2D human keypoints as reliable correspondences and 3D human meshes as robust structures to enhance bundle adjustment and produce coherent, metric-scale reconstructions.
Please refer to the supplementary material for an ablation study on camera/scene scale initialization and the impact of the number of cameras.
Qualitative Results: See Figs. 1 and 3 for qualitative results on EgoHumans and EgoExo4D and our Sup. Mat. for an in-the-wild demo with images we captured using a minimal setup of just two cell phones. Figure 3 shows intermediate optimization steps, where in the beginning, people are floating around in mid air. After joint optimization, their feet are consistent with the environment, without any explicit contact constraints. The structure also improves as pointmaps are more coherent around the ground. Please note that the human point cloud is not used as prior for human reconstruction, i.e. noisy point clouds or gaps do not directly affect the human reconstruction accuracy. Despite clear improvements in human localization, we sometimes observe scenes with slightly uneven ground planes indicating that scene reconstruction remains a challenging problem. Nonetheless, our approach improves camera metrics and scene reconstruction qualitatively (see Fig. S.3).